OpenStudy (anonymous):

what is the differentiation of cos x -xe^x, if x = 1.0 what is the solution

OpenStudy (anonymous):

sec - e

OpenStudy (anonymous):

is this the right answer

OpenStudy (anonymous):

\[f(x)=\cos x-xe ^{x}\] \[f \prime(x)=-\sin x-[(x)\prime*e ^{x}+x.e ^{x}]=-\sin x-e ^{x}-xe ^{x}\] if x=1 \[-\sin 1-e\] and if the equation is \[f(x)=\cos (x-xe ^{x})\] \[f \prime(x)=-\sin (x-xe ^{x})*(1-e ^{x}-x e^{x})\]

OpenStudy (unklerhaukus):

\[f(x)=\cos (x)-xe ^{x}\]\[f ^\prime(x)=-\sin (x)-e ^{x}-xe ^{x}\] \[f ^\prime(1.0)=-\sin (1.0)-e ^{1.0}-1.0e ^{1.0}\]\[\qquad\qquad=-\sin(1)-e-e\]\[\qquad\qquad=-\sin(1)-2e\]